OpenStudy (anonymous):

write 2 1/11 as an improper fraction a.21/11 b.13/11 c.22/11 d.23/11

OpenStudy (anonymous):

d

OpenStudy (lalaly):

\[\large{\frac{(11 \times 2) +1}{11}}= \frac{23}{11}\]